# What Determines the Maximum Mass of AGN-assisted Black Hole Mergers?

### v1

@@ 3.3.2 Impact of BH populations
The BH mass parameters naturally play a significant role in determining merger masses. As expected, increasing the maximum mass of pre-existing single BHs leads to a higher upper limit on merger mass, while the overall merger rate remains relatively unchanged. However, we observe that the initial mass function index $\beta_{\mathrm{IMF}}$ and the initial maximum single mass $M_{\mathrm{BH,ini,max}}$ exhibit little relevance to either the merger rate or the highest merger mass, except for the highest mass of the isotropic distribution. This suggests that increases in individual BH masses are diluted by frequent binary-single interactions and binary formation, which promote hierarchical mergers across multiple generations. In contrast, the isotropic distribution, which has fewer disk BHs, relies more heavily on the initial single BH mass to determine the highest merger mass.
Parameters related to the spatial distribution significantly affect the merger rate but have minimal impact on merger masses. These parameters, $\gamma_{\rho}$, $N_{\mathrm{BH},\mathrm{ini}}$ and $r_{\mathrm{BH},\mathrm{out}}$ determine the BH number density near the SMBH (Eq.8), influencing the merger rate without substantially altering the highest merger mass. The inclination distribution, particularly $\beta_{v}$ in the Gaussian model, governs the number of BHs that settle into the disk, serving as potential seeds for binary formation and binary-single interactions. All these parameters are at least relevant to the merger rate, and changing $\gamma_{\rho}$ becomes a more effective way to increase the number density of disk BHs for isotropic model.
The pre-existing binary fraction $f_{\mathrm{pre}}$ and the initial maximum separation $R_{\max}$ both have negligible influence on both the merger rate and the highest BH mass. These two parameters mainly determine the gen=1 BH remnants, which is a secondary effect compared to the higher generation binaries both to merger rate and highest mass. What’s more, the separation rapidly decreases due to gas dynamical friction hardening as BHs migrate into the disk, making the initial separation less important.
@@ 3.3.3 Impact of disk parameters
In our analysis of disk parameter effects, we examine the influence of the AGN disk size ($r_{\mathrm{disk},\mathrm{in}},r_{\mathrm{disk},\mathrm{out}}$), accretion rate at the outer disk boundary $\dot{M}_{\mathrm{out}}$, the angular momentum transport rate $m_{\rm{AM}}$, Shakura-Sunyaev $\alpha$-parameter $\alpha_{\mathrm{SS}}$, the conversion efficiency $\epsilon$, and the star-formation initial mass function index $\beta_{\mathrm{cre}}$. Some disk variables, like $m_{\rm{AM}}$, are analyzed beyond their typical range ($m_{\rm{AM}}$ is normally 0.1-0.2) to assess their influence. The AGN lifetime $t_{\mathrm{AGN}}$ is unique in that it only extends the simulation duration, thereby increasing the maximum merger mass without affecting the merger rate a lot, as discussed in Sec. 3.2.
@@
While all disk parameters influence disk properties, minor variations in these parameters may result in negligible changes in scale height or gas density. However, certain variations can significantly impact the merger rate, primarily due to their effect on type I/II migration of BHs within the disk. Changes in disk properties can lead to variations in migration speed by a factor of $\sim$ 2, thereby influencing the merger rate. Notably, parameters such as $m_{\rm{AM}}$ and $\epsilon$ induce by a factor of $\sim 2$ changes in migration speed at $r\sim 10^{-4}-10^{-1}$ pc. In contrast, the Shakura-Sunyaev viscosity parameter ($\alpha_{\rm{SS}}$) affects migration speed only at $r\lesssim 10^{-3}$ even when varied to an extremely small value $\alpha_{\rm{SS}}=0.005$, resulting in minimal impact on the merger rate, as illustrated in Fig. 13 and 14.
Modifying both the inner and outer sizes of the AGN disk significantly affects the maximum merger mass in the Gaussian distribution but generally has a minor impact on the isotropic distribution. With a migration speed variation of at most a factor of $<2$, we observe a noticeable decrease in the number of high-generation mergers when increasing either $r_{\mathrm{disk},\mathrm{out}}$ or $r_{\mathrm{disk},\mathrm{in}}$. This decline may be attributed to the substantial difference in number density at $r\sim 2\times 10^{-4}$ pc. However, this effect is prominent only in the Gaussian distribution, as the number of disk BHs in the isotropic model is insufficient to reveal a similar trend. Notably, when $\alpha_{\rm{SS}}$ takes very small values, it induces a similar trend in the maximum merger mass, it induces a similar trend in the maximum merger mass, but its effect is not significant enough to be considered a relevant parameter.
Surprisingly, direct changes to the accretion rate at outer radius $\dot{M}_{\mathrm{out}}$ have a less pronounced effect on the merger rate compared to variations in $m_{AM}$ and $\epsilon$. While increasing $\dot{M}_{\mathrm{out}}$ generally enhances the star formation rate, leading to minimal influence on the inner regions of the AGN disk. As a result, as long as $\dot{M}{\mathrm{out}}$ varies within its typical range, both the number density and migration speed remain largely unchanged, making its impact relatively insignificant. However, significant changes in $r_{\mathrm{disk},\mathrm{out}}$ could, in turn, influence the accretion rate at $r_{\mathrm{out}}=5\;\rm{pc}$, potentially altering the merger dynamics affecting the merger rate and mass.
@@ 3.3.4 Impact of SMBH mass
Changes in SMBH mass can lead to complex effects on both the merger rate and the maximum merger mass. As shown in Fig.13 and Fig.14, when the SMBH mass is below $10^{7}M_{\odot}$, the merger rate and maximum BH mass remain very low, primarily constrained by the initial number of BHs in the disk. As the SMBH mass increases, both the merger rate and maximum mass stabilize around $10^{6}-10^{7}M_{\odot}$.
@@
where we use the velocity dispersion estimate from Eq.127, gas dynamical friction interactions from Eq.51, the sound speed relation $c_{s}=v_{\mathrm{kep}}(h/r)$, and the disk size from Eq. 29. This suggests that a higher SMBH mass leads to a slightly lower disk BH number density, reducing the binary-single interaction rate and binary formation rate, ultimately decreasing the overall merger rate. This effect becomes significant when the disk BH population is insufficient for frequent hierarchical mergers, as seen in the isotropic distribution. However, for the Gaussian distribution, even with a factor of $310^{-1/4}\approx 1/$ 3 decrease in number density compared to fiducial mass, the number of BHs remains sufficient to sustain enough hierarchical mergers.
@@ 3.3.5 Implications
Combining the results from Sec. 3.2 with our previous discussion, we identify several key factors influencing both the merger rate and the highest merger mass. If a disk parameter modifies the Type I/II migration rate on the order of $\mathcal{O}(1)$, it typically affects the merger mass by facilitating the migration of binaries into high-density gas regions, thereby enhancing binary hardening. Additionally, if a disk parameter variation moves the location of a migration trap (due to discontinuities in disk properties) at $\mathcal{O}(1)$ scale, it can slow down inward migration at small radii, increasing the likelihood of hierarchical mergers and contributing to a higher maximum merger mass.
@@ 4.1 Convergence Test
We examined the impact of numerical resolution on the merger rate and the highest merger mass by varying the number of radial cells ($N_{\mathrm{cell}}$), the number of mass cell ($N_{\rm{mass}}$) and the time step parameter $\eta_{t}$. When adjusting the number of radial cells to $N_{\mathrm{cell}}=80,160$, the merger rate $\Gamma_{\rm{merger}}$ changed by only $\sim$2for both Gaussian model, while the top 1% mass $M_{\rm{top}}$ changes 15% and 8% for Gaussian model. In contrast, for the isotropic model, both the merger rate and top 1% mass both changed by less than 4%. The slower convergence of the Gaussian model likely stems from its higher merger rate, where 10,000 mergers may not be sufficient for full statistical convergence.
When varying the number of mass cells to $N_{\rm{mass}}=80,120$, the merger rate $\Gamma_{\rm{merger}}$ and the top 1% mass $M_{\rm{top}}$ exhibited only minor fluctuations of 1%-5% for both isotropic and Gaussian model. Additionally, we tested the time step parameter at $\eta_{t}=0.05$ and $\eta_{t}=0.2$. For both Gaussian and isotropic model, the merger rate and highest mass changes by $\sim$15for $\eta_{t}=0.05$ and $\eta_{t}=0.2$. However, with the $\eta_{t}=0.05$ applied, the time cost becomes incredibly large (increase by more than 5 times from $\eta_{t}=0.1$). This larger variation arises because the weak scattering diffusion term scales as $\sqrt{\Delta t}$, as indicated in Eq. 75. Although convergence is slower for the Gaussian distribution, our main conclusions regarding the trends of parameter variations remain unaffected by numerical resolution. In the next section, we provide potential methods to solve this issue.
4.2 Limitations
In our program, we adopt a highly simplified model at the statistical background level. We assume a collisionless BH disk with scale height $h_{\mathrm{DBH}}$, neglecting both binary-single interactions and binary formation processes. While this approximation may hold in non-collisional or weakly collisional systems, it introduces several limitations and potential inaccuracies in the context of AGN disks, where dynamical interactions are expected to be significant.
@@ 4.2 Limitations
Second, neglecting interactions leads to an overestimation of the local BH number density and thus an inflated binary formation rate. Although we apply a correction at the Monte Carlo stage using Eq. 70 to account for number reduction due to binary formation, this only considers binaries formed within the simulated BH sample. Background BHs that are not explicitly tracked can also form binaries and deplete the local population—an effect not included in our current model. As a result, the background BH number density remains systematically overestimated.
@@
Third, by ignoring binary formation in the background, we implicitly assume that all binary partners are drawn from the first-generation BH population. In other words, only 1g+Ng hierarchical mergers are included, while Mg+Ng mergers involving two higher-generation BHs are omitted. This constraint affects the merger weighting in Eqs. 133 and 134, leading to an overestimated merger rate and potentially underestimating the maximum BH mass attainable in the system.
@@
A potential solution to address all three issues is to explicitly track single and binary BH populations at the background level, while incorporating higher-generation mergers. However, this requires accurately modeling binary-single interactions, including ejection and re-entry processes, without significantly increasing computational cost—a challenge we aim to tackle in future work.
Finally, the artificially high BH number density may have additional consequences. At late times, the total BH mass could approach or even exceed the local AGN disk mass, leading to significant dynamical feedback on the disk itself. These concerns are model-dependent, but they highlight the importance of developing a more realistic treatment of BH dynamics in AGN disks.
@@ 4.3 Migration Traps and Extreme Mass-ratio Inspirals
Figure 16: The migration normalized factor $C_{\rm{I}}$ and $C_{\rm{I/II}}$ for type I and type I/II migration. The black line represent our simplified model $f_{\mathrm{mig}}=2$ using in the simulation. Additionally, we plot torque models from Jiménez & Masset (2017) (jim17) for both isothermal (jim17_iso, blue) and non-isothermal cases (jim17_tot, orange), Paardekooper et al. (2010) (p10, green) for the isothermal case, and Bellovary et al. (2016) (b16, purple), which adopts the formula from Paardekooper et al. (2010). The gray dotted line indicates regions where the torque results in outward migration.
As discussed in Section 2.4.3, a BH orbiting within a gas disk exchanges angular momentum with its surroundings, leading to net inward radial migration. Paardekooper & Mellema (2006) demonstrated that, under certain conditions, migration can proceed outward, resulting in the formation of a migration trap where outwardly migrating objects meet inwardly migrating ones. Such migration traps have been predicted in protoplanetary disks (Lyra et al., 2010) and in AGN disks, as suggested by Bellovary et al. (2016) using the torque formula from Paardekooper et al. (2010). Numerous studies highlight their significance in contributing to hierarchical mergers (McKernan et al., 2012; Yang et al., 2019; Tagawa et al., 2020b; Santini et al., 2023; Gilbaum & Stone, 2022; Vaccaro et al., 2024).
However, recent studies have challenged the existence of migration traps in AGN disks. Peng & Chen (2021) found that the outward migration region vanishes when considering the torque exerted by mini-disks surrounding black holes due to gas drag. Pan & Yang (2021) reached a similar conclusion, accounting for the influence of density wave generation and headwind effects. Grishin et al. (2024), employing an updated torque formula from Jiménez & Masset (2017), also observed the disappearance of the outward migration region. However, Grishin et al. (2024) introduced the role of thermal torque, generated by the thermal response of the AGN to black holes accreting via their own mini-disks, which in turn facilitates the formation of migration traps. Gangardt et al. (2024) then applied this mechanism to the AGN disk model from Thompson et al. (2005) and Sirko & Goodman (2003), successfully recovering migration traps at various radii.
The aforementioned studies consider only Type I migration torque, which is the direct summation of corotation torque and Lindblad torque. In AGN disks, the gap-opening condition is typically satisfied when $K\gtrsim 20$ (Kanagawa et al., 2018), where $K$ is defined in Eq. 38. Given that our disk model is very thin $h/r\sim 10^{-3}$, this condition is easily met at small radii. We thus normalize the type I and type I/II migration interaction rates as $\Gamma_{\rm{I}}=C_{\rm{I}}\Gamma_{0}$ and $\Gamma_{\rm{I/II}}=C_{\rm{I/II}}\Gamma_{0}/(1+0.04K)$, where
@@
and adopt the formulation from Kanagawa et al. (2018) to account the gap opening:
@@
$C_{\rm{L}}=\begin{cases}-2.5+0.1\nabla_{\Sigma}+0.5\nabla_{T}&\textit{p10\_iso}\\ (-2.5+0.1\nabla_{\Sigma}-1.7\nabla_{T})/\gamma_{\mathrm{gas}}&\textit{p10\_ad}\\ -2.34+0.1\nabla_{\Sigma}-1.5\nabla_{T}&\textit{jim17\_iso}\\ \left(-2.34+0.1\nabla_{\Sigma}-1.5\nabla_{T}\right)f_{\gamma}\left(\frac{\chi}{\chi_{c}}\right)&\textit{jim17\_tot}\\ \end{cases}$ (139)
@@
$C_{\rm{C}}=\begin{cases}1.65-1.1\nabla_{\Sigma}-1.4\nabla_{T}&\textit{p10\_iso}\\ (1.65-1.1\nabla_{\Sigma}+7.9\nabla_{S}/\gamma_{\mathrm{gas}})/\gamma_{\mathrm{gas}}&\textit{p10\_ad}\\ 0.976-0.64\nabla_{\Sigma}+\nabla_{T}&\textit{jim17\_iso}\\ (0.46-0.96\nabla_{\Sigma}+1.8\nabla_{T})/\gamma_{\mathrm{gas}}&\textit{jim17\_tot}\\ \end{cases}$ (140)

### v2

@@ 3.3.3 Impact of BH populations
The BH mass parameters naturally play a significant role in determining merger masses. As expected, increasing the maximum mass of preexisting single BHs leads to a higher upper limit on merger mass, while the overall merger rate remains relatively unchanged. However, we observe that the initial mass function index $\beta_{\mathrm{IMF}}$ and the initial maximum single mass $M_{\mathrm{BH,ini,max}}$ exhibit little relevance to either the merger rate or the highest merger mass, except for the highest mass of the isotropic distribution. This suggests that increases in individual BH masses are diluted by frequent binary-single interactions and binary formation, which promote hierarchical mergers across multiple generations. In contrast, the isotropic distribution, which has fewer disk BHs, relies more heavily on the initial single BH mass to determine the highest merger mass.
Parameters related to the spatial distribution significantly affect the merger rate but have minimal impact on merger masses. These parameters, $\gamma_{\rho}$, $N_{\mathrm{BH},\mathrm{ini}}$ and $r_{\mathrm{BH},\mathrm{out}}$ determine the BH number density near the SMBH (Eq.A8), influencing the merger rate without substantially altering the highest merger mass. The inclination distribution, particularly $\beta_{v}$ in the Gaussian model, governs the number of BHs that settle into the disk, serving as potential seeds for binary formation and binary-single interactions. All these parameters are at least relevant to the merger rate, and changing $\gamma_{\rho}$ becomes a more effective way to increase the number density of disk BHs for the isotropic model.
The preexisting binary fraction $f_{\mathrm{pre}}$ and the initial maximum separation $R_{\max}$ both have negligible influence on both the merger rate and the highest BH mass. These two parameters mainly determine the gen=1 BH remnants, which is a secondary effect compared to the higher generation binaries both to merger rate and highest mass. What’s more, the separation rapidly decreases due to gas dynamical friction hardening as BHs migrate into the disk, making the initial separation less important.
@@ 3.3.4 Impact of disk parameters
In our analysis of disk parameter effects, we examine the influence of the AGN disk size ($r_{\mathrm{disk},\mathrm{in}},r_{\mathrm{disk},\mathrm{out}}$), accretion rate at the outer disk boundary $\dot{M}_{\mathrm{out}}$, the angular momentum transport rate $m_{\rm{AM}}$, Shakura-Sunyaev $\alpha$-parameter $\alpha_{\mathrm{SS}}$, the conversion efficiency $\epsilon$, and the star-formation initial mass function index $\beta_{\mathrm{cre}}$. Some disk variables, like $m_{\rm{AM}}$, are analyzed beyond their typical range ($m_{\rm{AM}}$ is normally 0.1-0.2) to assess their influence.
@@
While all disk parameters influence disk properties, minor variations in these parameters may result in negligible changes in scale height or gas density. However, certain variations can significantly impact the merger rate, primarily due to their effect on type I/II migration of BHs within the disk. Changes in disk properties can lead to variations in migration speed by a factor of $\sim$ 2, thereby influencing the merger rate. Notably, parameters such as $m_{\rm{AM}}$ and $\epsilon$ induce by a factor of $\sim 2$ changes in migration speed at $r\sim 10^{-4}-10^{-1}$ pc. In contrast, the Shakura-Sunyaev viscosity parameter ($\alpha_{\rm{SS}}$) affects migration speed only at $r\lesssim 10^{-3}$ even when varied to an extremely small value $\alpha_{\rm{SS}}=0.005$, resulting in minimal impact on the merger rate, as illustrated in Fig. 8 and 9.
Modifying both the inner and outer sizes of the AGN disk significantly affects the maximum merger mass in the Gaussian distribution but generally has a minor impact on the isotropic distribution. With a migration speed variation of at most a factor of <2<2, we observe a noticeable decrease in the number of high-generation mergers when increasing either $r_{\mathrm{disk},\mathrm{out}}$ or $r_{\mathrm{disk},\mathrm{in}}$. This decline may be attributed to the substantial difference in number density at $r\sim 2\times 10^{-4}$ pc. However, this effect is prominent only in the Gaussian distribution, as the number of disk BHs in the isotropic model is insufficient to reveal a similar trend. Notably, when $\alpha_{\rm{SS}}$ takes very small values, it induces a similar trend in the maximum merger mass, it induces a similar trend in the maximum merger mass, but its effect is not significant enough to be considered a relevant parameter.
Surprisingly, direct changes to the accretion rate at outer radius $\dot{M}_{\mathrm{out}}$ have a less pronounced effect on the merger rate compared to variations in $m_{AM}$ and $\epsilon$, while increasing $\dot{M}_{\mathrm{out}}$ generally enhances the star-formation rate, leading to minimal influence on the inner regions of the AGN disk. As a result, as long as $\dot{M}_{\mathrm{out}}$ varies within its typical range, both the number density and migration speed remain largely unchanged, making its impact relatively insignificant. However, significant changes in $r_{\mathrm{disk},\mathrm{out}}$ could, in turn, influence the accretion rate at $r_{\mathrm{out}}=5\;\rm{pc}$, potentially altering the merger dynamics affecting the merger rate and mass.
@@ 3.3.5 Impact of SMBH mass
Changes in SMBH mass can lead to complex effects on both the merger rate and the maximum merger mass. As shown in Figs. 8 and 9, when the SMBH mass is below $10^{7}M_{\odot}$, the merger rate and maximum BH mass remain very low, primarily constrained by the initial number of BHs in the disk. As the SMBH mass increases, both the merger rate and maximum mass stabilize around $10^{6}-10^{7}M_{\odot}$.
@@
where we use the velocity dispersion estimate from Eq.F20, gas dynamical friction interactions from Eq.D16, the sound speed relation $c_{s}=v_{\mathrm{kep}}(h/r)$, and the disk size from Eq. B14. This suggests that a higher SMBH mass leads to a slightly lower disk BH number density, reducing the binary-single interaction rate and binary formation rate, ultimately decreasing the overall merger rate. This effect becomes significant when the disk BH population is insufficient for frequent hierarchical mergers, as seen in the isotropic distribution. However, for the Gaussian distribution, even with a factor of $10^{-1/4}\approx 1/3$ decrease in number density compared to fiducial mass, the number of BHs remains sufficient to sustain enough hierarchical mergers.
@@ 3.3.6 Implications
Combining the results from Sec. 3.2 with our previous discussion, we identify several key factors influencing both the merger rate and the highest merger mass. If a disk parameter modifies the type I/II migration rate on the order of $\mathcal{O}(1)$, it typically affects the merger mass by facilitating the migration of binaries into high-density gas regions, thereby enhancing binary hardening. Additionally, if a disk parameter variation moves the location of a migration trap (due to discontinuities in disk properties) at an $\mathcal{O}(1)$ scale, it can slow down inward migration at small radii, increasing the likelihood of hierarchical mergers and contributing to a higher maximum merger mass.
@@ 4.1 Convergence test
We examined the impact of numerical resolution on the merger rate and the highest merger mass by varying the number of radial cells ($N_{\mathrm{cell}}$), the number of mass cell ($N_{\rm{mass}}$) and the time step parameter $\eta_{t}$. When adjusting the number of radial cells to $N_{\mathrm{cell}}=80,160$, the merger rate $\Gamma_{\rm{merger}}$ changed by only $\sim$2% for both Gaussian models, while the top 1% mass $M_{\rm{top}}$ changes 15% and 8% for Gaussian models. In contrast, for the isotropic model, both the merger rate and top 1% mass both changed by less than 4%. The slower convergence of the Gaussian model likely stems from its higher merger rate, where 10,000 mergers may not be sufficient for full statistical convergence.
When varying the number of mass cells to $N_{\rm{mass}}=80,120$, the merger rate $\Gamma_{\rm{merger}}$ and the top 1% mass $M_{\rm{top}}$ exhibited only minor fluctuations of 1%-5% for both isotropic and Gaussian models. Additionally, we tested the time step parameter at $\eta_{t}=0.05$ and $\eta_{t}=0.2$. For both Gaussian and isotropic model, the merger rate and highest mass changes by $\sim$15% for $\eta_{t}=0.05$ and $\eta_{t}=0.2$. However, with the $\eta_{t}=0.05$ applied, the time cost becomes incredibly large (increase by more than 5 times from $\eta_{t}=0.1$). This larger variation arises because the weak scattering diffusion term scales as $\sqrt{\Delta t}$, as indicated in Eq. E11. Although convergence is slower for the Gaussian distribution, our main conclusions regarding the trends of parameter variations remain unaffected by numerical resolution. In the next section, we provide potential methods to solve this issue.
4.2 Limitations and neglected effects
In this work, we adopt a highly simplified model at the statistical background level. We assume a collisionless BH disk with scale height $h_{\mathrm{DBH}}$, neglecting both binary-single interactions and binary formation processes. While this approximation may hold in noncollisional or weakly collisional systems, it introduces several limitations and potential inaccuracies in the context of AGN disks, where dynamical interactions are expected to be significant.
@@ 4.2 Limitations and neglected effects
Second, neglecting interactions leads to an overestimation of the local BH number density and thus an inflated binary formation rate. Although we apply a correction at the Monte Carlo stage using Eq. E6 to account for number reduction due to binary formation, this only considers binaries formed within the simulated BH sample. Background BHs that are not explicitly tracked can also form binaries and deplete the local population—an effect not included in our current model. As a result, the background BH number density remains systematically overestimated.
@@
Third, by ignoring binary formation in the background, we implicitly assume that all binary partners are drawn from the first-generation BH population. In other words, only 1g+Ng hierarchical mergers are included, while Mg+Ng mergers involving two higher-generation BHs are omitted. This constraint affects the merger weighting in Eqs. 1 and 2, leading to an overestimated merger rate and potentially underestimating the maximum BH mass attainable in the system.
@@
A potential solution to address all three issues is to explicitly track single and binary BH populations at the background level, while incorporating higher-generation mergers. However, this requires accurately modeling binary-single interactions, including ejection and reentry processes, without significantly increasing computational cost—a challenge we aim to tackle in future work.
On addition to the modeling assumptions described above, our simulations neglect several physical processes. Notably, we do not account for the spin evolution of BHs. High-spin BHs can result in significant gravitational recoil velocities following mergers (Rezzolla et al., 2008). However, since the majority of mergers occur in close proximity to the central SMBH (as shown in Fig. 4), we do not expect the recoiling remnants to become unbound from the SMBH. While spin-induced differences in recoil velocity could slightly alter the return timescale of remnant BHs to the disk, such variations are likely minor due to the high ambient gas density. These effects are expected to have limited impact, especially considering that we focus primarily on the fastest mergers in the simulation.
Another major omission is the feedback from BHs. Radiative and kinetic feedback from accreting BHs can heat the surrounding gas or drive outflows, suppressing further accretion at late times (Sadowski et al., 2015; Jiang et al., 2014). Furthermore, our assumption of artificially high BH number densities may lead to unphysical scenarios in which the total BH mass approaches or exceeds the local AGN disk mass. In such cases, dynamical feedback could significantly alter the disk’s structure and stability. These effects are highly model-dependent, emphasizing the need for more realistic treatments of BH–gas interactions and feedback in AGN disks. A promising direction would be to adopt time-dependent AGN disk models, such as that proposed by Epstein-Martin et al. (2025).
Another potentially important effect is the Kozai–Lidov (KL) mechanism, which can induce large-amplitude oscillations in eccentricity and inclination in hierarchical triple systems (Kozai, 1962; Lidov, 1962). In dense, gas-rich AGN environments, KL oscillations can drive BH triples to high eccentricities, enhancing merger rates (Tagawa et al., 2020b; Dittmann et al., 2024; Mockler et al., 2023) and possibly contributing to the formation of the most massive BHs in the system.
We also cap BH accretion rates slightly above the Eddington limit. However, recent radiation-hydrodynamic simulations indicate that BHs embedded in dense, optically thick environments can accrete at super-Eddington rates—up to $\sim 10^{3}$ times the Eddington limit—due to photon trapping and anisotropic radiation (Jiang et al., 2014; Ogawa et al., 2017; Toyouchi et al., 2024). Ignoring such extreme accretion regimes may lead to significant underestimation of BH mass growth from gas accretion, especially in regimes where gas accretion dominates over mergers (see Sec.  3.3.2).
Finally, our results are highly sensitive to the assumed AGN disk model. Variations of treatment in turbulence and magnetic fields in the AGN disk can lead to different disk structures and local properties. These differences in turn affect BH migration rates (Secunda et al., 2019) and binary accretion processes (Chen & Lin, 2023), ultimately influencing both the merger rate and the maximum BH mass attainable.
@@ 4.3 Migration traps and extreme mass-ratio inspirals
Figure 11: The migration normalized factor $C_{\rm{I}}$ and $C_{\rm{I/II}}$ for type I and type I/II migration. The black line represent our simplified model $f_{\mathrm{mig}}=2$ used in the simulation. Additionally, we plot torque models from Jiménez & Masset (2017) (jim17) for both isothermal (jim17_iso, blue) and non-isothermal cases (jim17_tot, orange), Paardekooper et al. (2010) (p10, green) for the isothermal case, and Bellovary et al. (2016) (b16, purple), which adopts the formula from Paardekooper et al. (2010). The gray dotted line indicates regions where the torque results in outward migration.
As discussed in Appendix D.3, a BH orbiting within a gas disk exchanges angular momentum with its surroundings, leading to net inward radial migration. Paardekooper & Mellema (2006) demonstrated that, under certain conditions, migration can proceed outward, resulting in the formation of a migration trap where outwardly migrating objects meet inwardly migrating ones. Such migration traps have been predicted in protoplanetary disks (Lyra et al., 2010) and in AGN disks, as suggested by Bellovary et al. (2016) using the torque formula from Paardekooper et al. (2010). Numerous studies highlight their significance in contributing to hierarchical mergers (McKernan et al., 2012; Yang et al., 2019; Tagawa et al., 2020b; Santini et al., 2023; Gilbaum & Stone, 2022; Vaccaro et al., 2024).
However, recent studies have challenged the existence of migration traps in AGN disks. Peng & Chen (2021) found that the outward migration region vanishes when considering the torque exerted by mini disks surrounding black holes due to gas drag. Pan & Yang (2021) reached a similar conclusion, accounting for the influence of density wave generation and headwind effects. Grishin et al. (2024), employing an updated torque formula from Jiménez & Masset (2017), also observed the disappearance of the outward migration region. However, Grishin et al. (2024) introduced the role of thermal torque, generated by the thermal response of the AGN to black holes accreting via their own mini disks, which in turn facilitates the formation of migration traps. Gangardt et al. (2024) then applied this mechanism to the AGN disk model from Thompson et al. (2005) and Sirko & Goodman (2003), successfully recovering migration traps at various radii. More recently, Gilbaum et al. (2025) investigated various mechanisms by which BHs might escape from these migration traps and assessed their implications for BH mass growth.
The aforementioned studies consider only type I migration torque, which is the direct summation of corotation torque and Lindblad torque. In AGN disks, the gap-opening condition is typically satisfied when $K\gtrsim 20$ (Kanagawa et al., 2018), where $K$ is defined in Eq. D3. Given that our disk model is very thin $h/r\sim 10^{-3}$, this condition is easily met at small radii. We thus normalize the type I and type I/II migration interaction rates as $\Gamma_{\rm{I}}=C_{\rm{I}}\Gamma_{0}$ and $\Gamma_{\rm{I/II}}=C_{\rm{I/II}}\Gamma_{0}/(1+0.04K)$, where
@@
and adopt the formulation from Kanagawa et al. (2018) to account for the gap opening:
@@
[…]
@@
[…]